OpenStudy (anonymous):

Give the vector equation of the line passing through P = (1,-2) and Q = (3,0).

OpenStudy (perl):

the vector equation will have the form r(t) = P + t ( Q- P)

OpenStudy (perl):

r(t) = (1,-2) + t [ (3,0) - (1,-2) ]

OpenStudy (perl):

r(t) = (1,-2) + t [ ( 3 - 1 , 0 - -2 ) ] r(t) = (1,-2) + t ( 2, 2 )

OpenStudy (anonymous):

Thanks a bunch!